OpenStudy (yamyam70):

I need help putting up the equation A brother and sister have weight (pounds) that varies as the cube of height (feet) and they share the same proportionality constant. The sister is 6 feet tall and weighs 170 pounds. Her brother is 6'4" tall. How much does he wait?

OpenStudy (anonymous):

weight varies as the cube of height so basically weight = height^3

OpenStudy (anonymous):

and a proportionality constant :)

OpenStudy (yamyam70):

yes, what about the term , proportionality constant ?

OpenStudy (anonymous):

weight = constant * height^3

OpenStudy (anonymous):

Make units to how much an inch weights. Then compare those units to the boy.

OpenStudy (anonymous):

@HyperChemist 12 Inch = 1 foot ?

OpenStudy (yamyam70):

Yes

OpenStudy (anonymous):

Sorry phanta, you must of misunderstood something, which is weird because it's really clear. Divide the height of the person to the weight. Hence to make 1 inch equal whatever weight you get. Then, multiply how much one inch in height weights to the height of the boy.

OpenStudy (anonymous):

ok,so the sister is 6 x 12 = 60+12 = 72 Inches and the brother is 6 ft + 4 inch = 72 + 4 = 76 inches

OpenStudy (yamyam70):

That is indeed one way to solve the problem , but I'm looking for the solution where we use the equation , weight = k ( height)^3 then getting the weight of the boy.

OpenStudy (anonymous):

wait I'll re-write with k instand of "constant"

OpenStudy (yamyam70):

@phanta_seea , what is the equation for the boy's weight ?

OpenStudy (yamyam70):

:)

OpenStudy (anonymous):

we want to get the constant k to determine the boy's weight: weight = k * height^3 weight / k = height^3 k/ weight = 1/ height^3 <- reciprocal of the above k = weight/height^3

OpenStudy (anonymous):

@yamyam70 the equation for the boy's weight is just this: weight = k ( height)^3 however we need to know the proportionality constant to be able to gete the weight

OpenStudy (yamyam70):

weight =the value of k from girl times (6'4")^3 ?

OpenStudy (anonymous):

yes :)

OpenStudy (yamyam70):

@phanta_seea much thanks man :)

OpenStudy (anonymous):

original equation: weight = k * height^3 solve for k, by dividing height on both sides k = weight/height^3 plug in girl values: k = 170 pounds /(72 Inch)^3 with 6 ft = 72 Inch

OpenStudy (yamyam70):

@phanta_seea thanks :))

OpenStudy (anonymous):

no problem

OpenStudy (yamyam70):

something is wrong man

OpenStudy (yamyam70):

oh nevermind